# CorrSteer: Steering Improves Task Performance and Safety in LLMs through Correlation-based Sparse Autoencoder Feature Selection

### v1

@@ A.4 Ablation Study
Table 5: Pooling strategy comparison on Gemma 2 2B using CorrSteer-A. For single-token generation such as MMLU, MMLU-Pro, and BBQ, mean-pooling naturally achieves identical performance to max-pooling since only one token is generated, while all-token pooling shows degraded performance. Mean-pooling shows severe degradation on multi-token generation tasks, demonstrating the superiority of max-pooling.
Task Baseline Max-pooling Mean-pooling All-token
@@
To validate our design choice of using only positively correlated features, we conduct ablation experiments using negatively correlated features for steering. We compare two approaches: single-layer negative steering (CorrSteer-1 with negative features) and multi-layer negative steering (CorrSteer-A with negative features).
Table 6: Performance comparison between positive and negative correlation feature steering on Gemma 2 2B. Negative correlation features consistently show poor performance, validating our positive-only approach.
Task Baseline CorrSteer-A Negative-1 Negative-G
@@
The results demonstrate that negative correlation features provide minimal improvement in single-layer steering and often cause severe performance degradation in multi-layer steering. Notably, MMLU-Pro drops to 0.66% and BBQ Disambig to 12.15% with negative multi-layer steering, confirming that negative correlations often represent spurious patterns rather than causal relationships. Additionally, combining positive and negative features simultaneously yields inferior performance compared to positive-only selection. This validates our approach of using only positively correlated features, which aligns with the non-negative nature of SAE activations.
Table 7: Performance comparison between raw activation steering and SAE-decoded steering on Gemma 2 2B. Decoding adds SAE decoder bias term for the first layer, while Decoding-A adds multi-layer feature directions as CorrSteer-A.
Task Baseline Raw Activation Decoding-1 Decoding-A CorrSteer-A
@@
To validate the effectiveness of SAE-based sparse feature selection, we compare steering performance using raw residual stream activations. The results demonstrate a clear performance hierarchy: CorrSteer-A > SAE Decoding > Raw Activation across all evaluated tasks, which is explainable by Superposition Hypothesis (Elhage et al., 2022). One exception occurred in BBQ Disambig, where Decoding-1 shows better performance than CorrSteer-A. However, Decoding-1 failed to show robustness across benchmarks, frequently degrading performance while CorrSteer-A shows consistent performance across all tasks.
@@ A.5 Cross-Task Feature Transferability
Table 8: Cross-task feature transferability results on Gemma 2 2B. Features selected from source tasks (rows) are applied to target tasks (columns). Results show accuracy (%) with baseline performance in parentheses. MMLU-Pro results do not use constrained decoding, achieving 17.56% compared to unconstrained baseline (14.00%).
@@ A.6 Text Classification Validation
Mitigation Amplification
Category Baseline CorrSteer Biased CorrSteer
@@
Results demonstrate that correlation-selected features provide effective steering control across all demographic categories (Table 9). Our steering approach effectively reduces bias, with steered outputs showing substantially lower bias scores compared to biased baselines, supporting the generalizability of our approach across different domains and SAE training paradigms.
A demonstration of our bias mitigation results is available at https://huggingface.co/spaces/seonglae/CorrSteer, showcasing real-time steering capabilities.
A.7 Complete Feature Lists
@@ A.7.1 Gemma-2B
Figure: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with BBQ ambig on frequency in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with BBQ disambig on frequency in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with HarmBench on frequency in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with MMLU on coefficient in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with MMLU on frequency in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with MMLU-Pro on coefficient in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with MMLU-Pro on frequency in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with GSM8K on coefficient in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with GSM8K on frequency in each layer of Gemma 2 2B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with SimpleQA on coefficient in each layer of Gemma 2 2B.
@@
  • L4/680 references to standards, particularly in legal and medical contexts (coeff: 0.055, corr: 0.194)
  • L5/4460 references to legal cases and court rulings (coeff: 0.093, corr: 0.193)
  • L6/9777 expressions of agreement or dissent and the context surrounding them (coeff: 0.153, corr: 0.192)
  • L7/2431 phrases related to time management and constraints (coeff: 0.253, corr: 0.213)
  • L8/14209 HTML coding elements and formatting commands (coeff: 0.097, corr: 0.194)
  • L9/7856 terms related to penalties and scoring in sporting events (coeff: 0.257, corr: 0.196)
  • L10/2446 terms related to health and legal matters (coeff: 0.350, corr: 0.177)
  • L11/7954 legal terminology and references to court cases and proceedings (coeff: 0.164, corr: 0.194)
  • L12/1495 phrases related to the duration and continuity of experiences over time (coeff: 0.208, corr: 0.195)
  • L13/12119 references to various parameters and aspects within scientific or technical contexts (coeff: 0.148, corr: 0.194)
  • L14/6355 references to India and its cultural context (coeff: 0.377, corr: 0.201)
  • L15/4385 programming constructs related to class and method declarations in Java (coeff: 0.177, corr: 0.194)
  • L16/7182 expressions of enthusiasm or amazement (coeff: 0.457, corr: 0.216)
  • L17/6346 references to offices or organizational structures (coeff: 0.274, corr: 0.194)
  • L18/5258 terms related to legal charges and prosecutions (coeff: 0.843, corr: 0.200)
  • L19/4202 technical terms and concepts related to physical phenomena and their mathematical descriptions (coeff: 0.262, corr: 0.194)
  • L20/6557 legal terms and phrases related to criminal charges and legal proceedings (coeff: 0.953, corr: 0.220)
  • L21/13830 references to political leaders and government roles (coeff: 8.090, corr: 0.196)
  • L22/15897 phrases related to international relations and cooperation, particularly in the context of political statements and actions (coeff: 0.648, corr: 0.237)
  • L23/15190 terms related to health, well-being, and interventions for obesity and mental illness (coeff: 0.832, corr: 0.237)
  • L24/15228 references to political figures and their actions (coeff: 2.043, corr: 0.222)
  • L25/2531 expressions of political opinion regarding government spending and fiscal policies (coeff: 1.664, corr: 0.200)
Figure: Top correlated features with SimpleQA on frequency in each layer of Gemma 2 2B.
XSTest
Figure: Top correlated features with selected features from CorrSteer-P with XSTest on coefficient in each layer of Gemma 2 2B.
  • L1/4509 terms and concepts related to scientific and mathematical structures and functions (coeff: 1.940, corr: 0.333)
  • L2/4679 financial metrics and forecasts related to stock performance (coeff: 1.584, corr: 0.301)
  • L3/1326 legal terminology and references to statutes and claims (coeff: 4.088, corr: 0.256)
  • L4/12152 references to geographical locations and their associated attributes (coeff: 0.961, corr: 0.259)
  • L5/5939 terms related to signals and their coding in biological contexts (coeff: 3.391, corr: 0.248)
  • L6/4376 numerical values and specific formatting related to data structures or coding (coeff: 6.713, corr: 0.259)
  • L7/4886 representations of numerical data, particularly in scientific contexts (coeff: 3.074, corr: 0.286)
@@
Figure: Top correlated features with XSTest on frequency in each layer of Gemma 2 2B.
A.7.2 Llama-3.1-8B
@@
Figure: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with BBQ ambig on frequency in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with BBQ disambig on frequency in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with HarmBench on frequency in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with MMLU ambig on coefficient in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with MMLU on frequency in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with selected features from CorrSteer-P with MMLU-Pro ambig on coefficient in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with MMLU-Pro on frequency in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with SimpleQA on frequency in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with SimpleQA on frequency in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with XSTest on frequency in each layer of Llama 3.1 8B.
@@
Figure: Top correlated features with XSTest on frequency in each layer of Llama 3.1 8B.

### v2

@@ A.4 Ablation Study
Table 5: Pooling strategy comparison on Gemma-2 2B using CorrSteer-A. For single-token generation such as MMLU, MMLU-Pro, and BBQ, mean-pooling naturally achieves identical performance to max-pooling since only one token is generated, while all-token pooling shows degraded performance. Mean-pooling shows severe degradation on multi-token generation tasks, demonstrating the superiority of max-pooling.
Task Non-steered Max-pooling Mean-pooling All-token
@@
To validate our design choice of using only positively correlated features, we conduct ablation experiments using negatively correlated features for steering. We compare two approaches: single-layer negative steering (Corrsteer-S with negative features) and multi-layer negative steering (CorrSteer-A with negative features).
For Gemma-2 2B, negatively correlated features are bias- and choice-related:
  • L8/8123 questions asking for correctness of options (coeff: 3.725, corr: -0.133)
  • L17/9134 choice-related phrases and preferences (coeff: 2.379, corr: -0.451)
  • L19/15745 decision-making and choice expressions in social contexts (coeff: 9.740, corr: -0.464)
Table 6: Performance comparison between positive and negative correlation feature steering on Gemma-2 2B. Negative correlation features consistently show poor performance, validating our positive-only approach.
Task Non-steered CorrSteer-A Negative-S Negative-G
@@
The results demonstrate that negative correlation features provide minimal improvement in single-layer steering and often cause severe performance degradation in multi-layer steering. Notably, MMLU-Pro drops to 0.66% and BBQ Disambig to 12.15% with negative multi-layer steering, confirming that negative correlations often represent spurious patterns rather than causal relationships, as negative activations are often unrelated noise (Joseph Bloom, 2024). Additionally, combining positive and negative features simultaneously yields inferior performance compared to positive-only selection. This validates our approach of using only positively correlated features, which aligns with the non-negative nature of SAE activations.
Table 7: Performance comparison between raw activation steering and SAE-decoded steering on Gemma-2 2B. Decoding adds SAE decoder bias term for the first layer, while Decoding-A adds multi-layer feature directions as CorrSteer-A.
Task Non-steered Raw Activation Decoding-S Decoding-A CorrSteer-A
@@
To validate the effectiveness of SAE-based sparse feature selection, we compare steering performance using raw residual stream activations. The results demonstrate a clear performance hierarchy: CorrSteer-A > SAE Decoding > Raw Activation across all evaluated tasks, which is explainable by Superposition Hypothesis (Elhage et al., 2022). One exception occurred in BBQ Disambig, where Decoding-S shows better performance than CorrSteer-A. However, Decoding-S failed to show robustness across benchmarks, frequently degrading performance while CorrSteer-A shows consistent performance across all tasks.
@@ A.5 Cross-Task Feature Transferability
Table 8: Cross-task feature transferability results on Gemma-2 2B. Features selected from source tasks (rows) are applied to target tasks (columns). Results show accuracy (%) with non-steered model performance in parentheses. MMLU-Pro results do not use constrained decoding, achieving 17.56% compared to unconstrained non-steered model (14.00%).
@@ A.6 Text Classification Validation
Mitigation (Fairness $\uparrow$) Amplification (Bias $\uparrow$)
Category Non-steered CorrSteer Biased CorrSteer
@@
Results demonstrate that correlation-selected features provide effective steering control across all demographic categories (Table 9). For mitigation, CorrSteer surpasses the non-steered model across categories by improving fairness scores. For amplification, CorrSteer generally increases bias relative to the biased non-steered model, with the LGBTQ+ row as an exception to be audited. A demonstration of our bias mitigation results is available at https://huggingface.co/spaces/seonglae/CorrSteer, showcasing real-time steering capabilities.
A.7 Framework Implications
CorrSteer leverages generation-time activations for multi-token, multi-layer SAE-based steering, and our experiments are enabled by Gemma Scope (Lieberum et al., 2024) and LLaMA Scope (He et al., 2024), the only open releases providing SAEs across all residual stream layers.
The proposed framework demonstrates the practical utility of SAE in real-world LLM inference, addressing critical concerns such as safe reasoning, bias mitigation, and resistance to jailbreaking. This research demonstrates that SAE-based control mechanisms offer a promising direction for both understanding and improving LLM behavior. The framework’s ability to operate through an interpretable interface while maintaining or improving model performance suggests a concrete path toward safer, more transparent AI.
A.8 Complete Feature Lists
@@ A.8.1 Gemma-2B
Figure 8: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of Gemma-2 2B.
@@
Figure 9: Top correlated features with BBQ ambig on frequency in each layer of Gemma-2 2B.
@@
Figure 10: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of Gemma-2 2B.
@@
Figure 11: Top correlated features with BBQ disambig on frequency in each layer of Gemma-2 2B.
@@
Figure 12: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of Gemma-2 2B.
@@
Figure 13: Top correlated features with HarmBench on frequency in each layer of Gemma-2 2B.
@@
Figure 14: Top correlated features with selected features from CorrSteer-P with MMLU on coefficient in each layer of Gemma-2 2B.
@@
Figure 15: Top correlated features with MMLU on frequency in each layer of Gemma-2 2B.
@@
Figure 16: Top correlated features with selected features from CorrSteer-P with MMLU-Pro on coefficient in each layer of Gemma-2 2B.
@@
Figure 17: Top correlated features with MMLU-Pro on frequency in each layer of Gemma-2 2B.
@@
Figure 18: Top correlated features with selected features from CorrSteer-P with GSM8K on coefficient in each layer of Gemma-2 2B.
@@
Figure 19: Top correlated features with GSM8K on frequency in each layer of Gemma-2 2B.
@@
Figure 20: Top correlated features with selected features from CorrSteer-P with SimpleQA on coefficient in each layer of Gemma-2 2B.
@@
Figure 21: Top correlated features with XSTest on frequency in each layer of Gemma-2 2B.
A.8.2 LLaMA-3.1-8B
@@
Figure 22: Top correlated features with selected features from CorrSteer-P with BBQ ambig on coefficient in each layer of LLaMA-3.1 8B.
@@
Figure 23: Top correlated features with BBQ ambig on frequency in each layer of LLaMA-3.1 8B.
@@
Figure 24: Top correlated features with selected features from CorrSteer-P with BBQ disambig on coefficient in each layer of LLaMA-3.1 8B.
@@
Figure 25: Top correlated features with BBQ disambig on frequency in each layer of LLaMA-3.1 8B.
@@
Figure 26: Top correlated features with selected features from CorrSteer-P with HarmBench on coefficient in each layer of LLaMA-3.1 8B.
@@
Figure 27: Top correlated features with HarmBench on frequency in each layer of LLaMA-3.1 8B.
@@
Figure 28: Top correlated features with selected features from CorrSteer-P with MMLU ambig on coefficient in each layer of LLaMA-3.1 8B.
@@
Figure 29: Top correlated features with MMLU on frequency in each layer of LLaMA-3.1 8B.
@@
Figure 30: Top correlated features with selected features from CorrSteer-P with MMLU-Pro ambig on coefficient in each layer of LLaMA-3.1 8B.
@@
Figure 31: Top correlated features with MMLU-Pro on frequency in each layer of LLaMA-3.1 8B.
@@
Figure 32: Top correlated features with SimpleQA on frequency in each layer of LLaMA-3.1 8B.
@@
Figure 33: Top correlated features with SimpleQA on frequency in each layer of LLaMA-3.1 8B.
@@
Figure 34: Top correlated features with XSTest on frequency in each layer of LLaMA-3.1 8B.
@@
Figure 35: Top correlated features with XSTest on frequency in each layer of LLaMA-3.1 8B.